OpenStudy (anonymous):

The ordered pairs (1, 1), (2, 4), (3, 9), (4, 16), and (5, 25) represent a function. What is a rule that represents this function?

OpenStudy (agent0smith):

This may help... \[\Large 1^2 = 1\]\[\Large 2^2 = 4\]\[\Large 3^2 = 9\]\[\Large 4^2 = 16\]\[\Large 5^2 = 25\]

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

would it be y = x squared?

OpenStudy (agent0smith):

That is correct :)

OpenStudy (anonymous):

thank you(:

OpenStudy (agent0smith):

Welcome :) if you continued what i wrote above, you'd get the rule you found... \[\Large 5^2 = 25\]\[\Large 6^2 = 36 \] or \[\Large x^2 = y\]